The engine of a car moving at a speed of 30 km / h, develops a traction of 2700 N. Find the engine power.


1
Expert's answer
2021-04-19T17:08:50-0400

The power is the product of the developed force "F = 2700N" and the moving velocity "v = 30km/h = \frac{75}{9}m/s" . Thus:


"P = Fv = 2700N\cdot \dfrac{75}{9}m/s = 22500\space W"

Answer. 22500 W.